# On computability of equilibrium states

### v2

@@ 1. Introduction
    There is no computable $\widehat{T}$-invariant probability measure.
Strategy of the proof. We discuss below the strategies applied in the proofs of our main results. Let us first concentrate on the proof of Theorem A. Here is a precise version of Theorem A.
@@
There exists an algorithm that satisfies the following property: For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a Borel probability measure $\mu_{n}\in\mathcal{P}(X)$ supported on finitely many points of $\mathcal{S}$ such that the Wasserstein–Kantorovich distance
@@
after inputting the following data in this algorithm:
@@
    an algorithm outputting a net $G\subseteq\mathcal{S}$ of $X$ with any given precision (in the sense of Definition 3.16),
@@
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6),
@@
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
@@
The Wasserstein–Kantorovich metric $W_{\rho}$ on the set $\mathcal{P}(X)$ of Borel probability measures is defined in Section 2.1. Our proof of Theorem 1.1 is inspired by the proof of Theorem A in [BBRY11]. However, since we work on equilibrium states, additional methods are required. The most crucial of them is the technique of cones. We use it in Subsection 4.1 to estimate the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator (see the definition in Subsection 2.3) with explicit expressions for related constants (see Theorem 4.1). In the setting of smooth (or more generally Markov) expanding maps with smooth observable, related bounds were investigated by Ferrero ([Fe81]), Rychlik ([Ry89]), and Hunt ([Hu96]). In two groundbreaking papers [Liv95a, Liv95b] in 1995, Liverani used the technique of cones (based on an earlier work of Birkhoff [Bi57]) and gave a new upper bound of the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator for general (non-Markov) non-smooth hyperbolic systems with discontinuities. A critical ingredient of this method is the existence of a convex cone of functions that is mapped strictly inside itself by the Ruelle–Perron–Frobenius operator. Moreover, the diameter of the image with respect to some projective metric gives an estimate for the upper bound of the convergence rate (see Proposition 4.3). In the setting of smooth expanding maps on compact connected manifolds, Viana [Vi97, Chapter 2] provides a comprehensive estimate for the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator in Hölder norm. Notably, because of the connectedness, some crucial local estimates of the oscillations of the functions in the cone can be extended globally (to the whole manifold). This, however, cannot work in our setting, since we allow for general disconnected spaces. To remedy this, we add a new parameter into the definition of cones used in [Vi97], effectively restricting to smaller cones of functions (see Definition 4.4), in order to estimate the diameters (see Proposition 4.6). This extra parameter takes care of the required global bounds on oscillations. Moreover, to make sure that the operator is a strict contraction on these new cones, we use a different version of the normalized Ruelle–Perron–Frobenius operator and utilize additional techniques (see Proposition 4.5). Finally, we use the convergence rate of the new operator to obtain the desired bounds for the normalized Ruelle–Perron–Frobenius operator. We need this new finer estimate for the convergence rate (Theorem 4.1) to establish the computability of the eigenfunction $u_{\varphi}$ of the Ruelle–Perron–Frobenius operator. Furthermore, in Subsection 4.2, we deduce the computability of the Jacobian of the equilibrium state whose explicit expression is given by
@@
In Subsection 4.3, we establish the lower-computable openness of some subsets $U$ and $W$ of $\mathcal{P}(X)$ by expressing them as unions of uniformly computable sequences of lower-computable open sets. It is worth noting that the lower-computable openness of $U$ depends on the computability of $T$, while the lower-computable openness of $W$ depends on the computability of the Jacobian function $J_{\varphi}(x)$ in addition to the computability of $T$. In Subsection 4.4, we complete the proof of Theorem 1.1 and apply it to establish Theorem B. Here is a precise version of Theorem B.
@@
There exists an algorithm that satisfies the following property: For each septet $(X,\,D,\,\rho,\,\mathcal{S},\,\mathcal{S}^{\prime},\,\varphi,\,T)$ satisfying the Assumption B in Section 1, this algorithm outputs $\mu_{n}\in\mathcal{P}(X)$ supported on finitely many points of $\mathcal{S}$ such that the Wasserstein–Kantorovich distance
@@
after inputting the following data in this algorithm:
@@
    two rational constants $a_{0}$ and $v_{0}$ satisfying that $\varphi$ is Hölder continuous with exponent $v_{0}$ and constant $a_{0}$ (in the sense of Definition 2.6),
@@
    three rational constants $\eta$, $\lambda,$ and $\xi$ satisfying that $T$ is distance-expanding with these constants (in the sense of Definition 2.1),
@@
Next, in Section 5, we discuss the strategy of our proof of Theorem C. By Theorem 1.2, it suffices to check that the septet (^$\bigl{(}\widehat{\mathbb{C}},\,\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w),\,t\varphi_{f},\,f\bigr{)}$ satisfies the Assumption B in Section 1, where $\mathbb{Q}^{2}$ is defined as $\{a+bi:a,\,b\in\mathbb{Q}\}$, and $w$ is a repelling periodic point of the rational map $f$. The main difficulty here is to show that $\mathcal{J}_{f}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$. We use the computability of the hyperbolic Julia sets proved by Braverman [Br04] to establish this statement. All of our analyses for Theorems 1.1 and 1.2 should still work after necessary modifications if we replace the algorithms in the inputs by oracles. For the simplicity of the presentation, we leave the details to the reader. Structure of the paper. In Section 2, we introduce our notations and review some basic notions and results in ergodic theory and thermodynamic formalism in particular. Section 3 covers some basics of Computable Analysis. Section 4 is dedicated to the proofs of Theorems 1.1 and 1.2. In Subsection 4.1, we prove that the convergence of the iterations of a normalized Ruelle–Perron–Frobenius operator is at an exponential rate with explicit expressions for the related constants. Based on this, we demonstrate the computability of topological pressure $P(T,\varphi)$ and the function $u_{\varphi}$, which leads to the computability of the Jacobian of the equilibrium state in Subsection 4.2. Then, we prove Lemma 4.13 in Subsection 4.3 and establish Theorems 1.1 and 1.2 in Subsection 4.4. In Section 5, we introduce some results on the dynamical and algorithmic aspects of hyperbolic rational maps and prove Theorem C. In Section 6, we establish Theorem E, from which we derive Theorem D. In Section 7, we investigate and construct some computable systems whose equilibrium states are all non-computable, establishing Theorems F and G. In Subsection 7.1, we recall that the computable map $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ given in [GHR10, Subsection 4.1] satisfies statements in Theorem F. Finally, in Subsection 7.2, we construct $\widehat{T}\coloneqq T\times T_{d}$ to establish Theorem G, where $T_{d}(x)\coloneqq dx\pmod{1}$ for each $d\in\mathbb{N}$ and each $x\in\mathbb{S}^{1}$.
@@ Acknowledgments
I. Binder was partially supported by an NSERC Discovery grant. Z. Li and Q. He were partially supported by NSFC Nos. 12101017, 12090010, 12090015, and BJNSF No. 1214021. Q. He was also supported by Peking University Funding Nos. 7101303303 and 6201001846. Y. Zhang was partially supported by NSFC Nos. 12161141002, 11871262, and 12271432.
@@ 2. Preliminaries on thermodynamic formalism
In this section, we go over some notations, key concepts, and useful results in ergodic theory and thermodynamic formalism in particular.
@@ 2.1. Notations
Denote {1, 2, $\mathbb{N}\coloneqq\{1,\,2,\,3,\,\ldots\}$ and $\mathbb{N}_{0}\coloneqq\{0\}\cup\mathbb{N}$. For each $k\in\mathbb{N}$, we fixed a bijection between $\mathbb{N}^{k}$ and $\mathbb{N}$ call it the “canonical bijection between $\mathbb{N}^{k}$ and $\mathbb{N}$”. Set *$\mathbb{N}^{*}\coloneqq\bigcup_{k\in\mathbb{N}}\mathbb{N}^{k}$, and fix a bijection between $\mathbb{N}^{*}$ and $\mathbb{N}$ that we call it the “canonical bijection between $\mathbb{N}^{*}$ and $\mathbb{N}$”. For a subset $A\subseteq X$, the characteristic function of $A$ is denoted by $\mathbbm{1}_{A}$. We write $\mathbbm{1}\coloneqq\mathbbm{1}_{X}$. Let $(X,\rho)$ be a compact metric space. Let $C(X)$ and $C^{*}(X)$ denote the Banach space of continuous functions from $X$ to $\mathbb{R}$ equipped with the norm $\|u\|_{\infty}\coloneqq\sup\{\lvert u(x)\rvert:x\in X\}$ and the dual space of $C(X)$, respectively. Given a constant $c>0$, a function $g\in C(X)$ is $c$-Lipschitz if $\lvert g(x)-g(y)\rvert\leqslant c\rho(x,y)$ for each pair of $x,\,y\in X$. We use $c\operatorname{-Lip}(X)$ to denote the set of real-valued $c$-Lipschitz continuous functions, and $B(x,r)=B_{\rho}(x,r)$ to denote the open ball of radius $r>0$ centered at $x\in X$.
Let $T\colon X\rightarrow X$ be a continuous map. The non-wandering set for $T$ is denoted by
@@
For all $\varphi\in C(X)$, $n\in\mathbb{N}$, and $x\in X$, denote
$S_{n}\varphi(x)\coloneqq\sum\limits_{k=0}^{n-1}\varphi\circ T^{k}(x).$
Let $\mathcal{P}(X)$ and $\mathcal{M}_{T}(X)$ denote the set of Borel probability measures and the set of $T$-invariant Borel probability measures on $X$, respectively. The Wasserstein–Kantorovich metric $W_{\rho}$ on $\mathcal{P}(X)$ is defined by
(2.2) $W_{\rho}(\mu,\nu)\coloneqq\sup\biggl{\{}\biggl{\lvert}\int\!f\,\mathrm{d}\mu-\int\!f\,\mathrm{d}\nu\biggr{\rvert}:f\in 1\operatorname{-Lip}(X)\biggr{\}}.$
@@ 2.2. Basic concepts in ergodic theory
We begin with a brief introduction of distance-expanding maps and topologically transitive (resp. exact) maps. See for example, [PU10, Chapters 3 and 4] for more details. Given constants $\eta>0$ and $\lambda>1$, assume that $(X,\rho)$ is a compact metric space and $T\colon X\rightarrow X$ is a continuous open map satisfying
@@
Then there exists a constant $\xi>0$ such that $B(T(x),\xi)\subseteq T(B(x,\eta))$ for each $x\in X$. Hence, for each $x\in X$ and each $n\in\mathbb{N}$, one can define the branches of the inverse map $T_{x}^{-n}\colon B(T^{n}(x),\xi)\rightarrow B(x,\eta)$ of $T^{n}$ in the sense of
$T^{-n}(A)=\bigcup_{x\in T^{-n}(y)}T^{-n}_{x}(A)$
for all $y\in X$ and $A\subseteq B(y,\xi)$ such that the following inequalities hold (see for example, [PU10, Section 4.1]):
(2.4) $\rho(T_{x}^{-n}(y),T_{x}^{-n}(z))\leqslant\lambda^{-n}\rho(y,z)\text{ for all }n\in\mathbb{N},\,x\in X,\,\text{and }y,\,z\in B(T^{n}(x),\xi).$
@@
Let $(X,\rho)$ be a compact metric space. Consider $\eta>0$, $\lambda>1$, and $\xi>0$. A map $T\colon X\rightarrow X$ is said to be distance-expanding with respect to the metric $\rho$ with constants $\eta$, $\lambda$, and $\xi$ if (2.3) and (2.4) both hold.
In a topological space $X$, recall that a continuous map $T\colon X\rightarrow X$ is topologically transitive if, for each pair of non-empty open sets $U,\,V\subseteq X$, there exists $n\in\mathbb{N}$ such that $T^{n}(U)\cap V\neq\emptyset$. Moreover, as a stronger property, we say a continuous map $T\colon X\rightarrow X$ is topologically exact if, for each non-empty open sets $U\subseteq X$, there exists $n\in\mathbb{N}$ such that $T^{n}(U)=X$. Under the above conventions, given constants $\eta>0$, $\lambda>1$, and $\xi>0$, a list of assumptions that are applied in the rest of this section is displayed below:
@@
    The map $T\colon X\rightarrow X$ is continuous, topologically transitive, open, and distance-expanding with respect to the metric $\rho$ constants $\eta$, $\lambda$, and $\xi$.
Next, we recall some notations and results in ergodic theory. For more details, we refer to [PU10, Chapter 2]. Let $X$ be a topological space. A measurable partition $\mathcal{A}$ of $X$ is a cover $\mathcal{A}=\{A_{j}:j\in J\}$ of $X$ consisting of finite or countably many mutually disjoint Borel sets. Unless otherwise stated, we assume that our partitions are all measurable. For each partition $\mathcal{A}$ of $X$ and $x\in X$, the unique element of $\mathcal{A}$ containing $x$ is denoted by $\mathcal{A}(x)$. Let $\mathcal{A}=\{A_{j}:j\in J\}$ and =$\mathcal{B}=\{B_{k}:k\in K\}$ be two covers of $X$, respectively, where $J$ and $K$ are corresponding index sets. We say that $\mathcal{A}$ is a refinement of $\mathcal{B}$ if for each $j\in J$, there exists $k=k(j)\in K$ such that $A_{j}\subseteq B_{k}$. The common refinement $\mathcal{A}\vee\mathcal{B}$ of $\mathcal{A}$ and $\mathcal{B}$ is a cover given by $\mathcal{A}\vee\mathcal{B}\coloneqq\{A_{j}\cap B_{k}:j\in J,\,k\in K\}.$ For each continuous map $T\colon X\rightarrow X$, by $T^{-1}(\mathcal{A})$ we denote the partition $\{T^{-1}(A_{j}):j\in J\}$, and for each $n\in\mathbb{N}$, write
@@
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous map, and $\varphi\colon X\rightarrow\mathbb{R}$ be a continuous function (which is often called a potential). Denote by $\mathcal{Y}$ the set of the sequences $\{\mathcal{V}_{n}\}_{n\in\mathbb{N}}$ of open finite covers of $X$ satisfying that the meshes of the covers $\mathcal{V}_{n}$ converge to zero as $n\rightarrow+\infty$, that is,
$\mathcal{Y}\coloneqq\bigl{\{}\{\mathcal{V}_{n}\}_{n\in\mathbb{N}}:\lim\limits_{n\to+\infty}\operatorname{mesh}(\mathcal{V}_{n})=0\bigr{\}},$
where the mesh of a cover $\mathcal{V}$ of $X$ is given by
$\operatorname{mesh}(\mathcal{V})\coloneqq\sup\{\operatorname{diam}(V):V\in\mathcal{V}\}.$
Define the topological pressure as the supremum of the following limits taken over all sequences $\{\mathcal{V}_{n}\}\in\mathcal{Y}$:

### v3

@@ 1. Introduction
    There is no computable $\widehat{T}-$invariant Borel probability measure.
Strategy of the proofs.
In the following, we discuss the strategies applied to the proofs of our main results.
Let us first concentrate on the proof of Theorem A. Here is a detailed version of Theorem A.
@@
There exists an algorithm with the following property:
For each quintet $(X,\,\rho,\,\mathcal{S},\,\varphi,\,T)$ satisfying the Assumption A in Section 1, this algorithm outputs a Borel probability measure $\mu_{n}\in\mathcal{P}(X)$ supported on finitely many points of $\mathcal{S}$ such that the Wasserstein–Kantorovich distance
@@
where the measure $\mu$ is the unique equilibrium state for $T$ and $\varphi$, after inputting the following data:
@@
    an algorithm outputting a net $G\subseteq\mathcal{S}$ of $X$ with any precision (in the sense of Definition 3.16),
@@
    two rational constants $a_{0}>0$ and $v_{0}\in(0,1]$ with $\varphi\in C^{0,v_{0}}(X)$ and $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$,
@@
    three rational constants $\eta>0$, $\lambda>1,$ and $\xi>0$ satisfying that $T$ is distance-expanding with constants $\eta,\,\lambda$, and $\xi$ (in the sense of Definition 2.1),
@@
The result of Theorem 1.1 is uniform, in the sense that there is a single algorithm that takes the associated algorithms and constants as parameters and computes the corresponding equilibrium state.
Our proof of Theorem 1.1 adapts some ideas from the proof of Theorem A in [BBRY11]. In addition, since we work on equilibrium states, the Jacobians of the desired measures become the focus of investigation. To establish the computability of the Jacobians, we need to estimate the convergence rate of the iterations of normalized Ruelle–Perron–Frobenius operators (see the definition in Subsection 2.3) with explicit expressions for related constants (see Theorem 4.1).
To estimate the rate, we introduce the cone technique into the study of Computable Analysis by inventing a new class of cones and designing a novel approach to using cones in thermodynamic formalism in order to control the constants for the purpose of Computable Analysis.
@@
The cone technique is an important tool for estimating the convergence rate of the iterations of normalized Ruelle–Perron–Frobenius operators. This technique, rooted in projective geometry, allows us to bound the contraction properties of transfer operators by restricting functions to some special convex subsets (cones). In the setting of smooth (or more generally, Markov) expanding maps with smooth observable, related bounds were investigated by Ferrero ([Fe81]), Rychlik ([Ry89]), and Hunt ([Hu96]). In two groundbreaking papers [Liv95a, Liv95b] in 1995, Liverani used the cone technique (based on an earlier work of Birkhoff [Bi57]) and gave a new upper bound of the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator for general (non-Markov) non-smooth hyperbolic systems with discontinuities. A critical ingredient of this method is the existence of a convex cone of functions that is mapped strictly inside itself by the Ruelle–Perron–Frobenius operator. Moreover, the diameter of the image with respect to some projective metric gives an estimate for the upper bound of the convergence rate (see Proposition 4.3).
In the setting of smooth expanding maps on compact connected manifolds, Viana [Vi97, Chapter 2] provides a comprehensive estimate for the convergence rate of the iterations of a normalized Ruelle–Perron–Frobenius operator in Hölder norm. Notably, because of the connectedness, some crucial local estimates of the oscillations of the functions in the cone can be extended globally (to the whole manifold). This, however, cannot work in our setting, since we allow for general disconnected spaces. To remedy this, we add a new parameter into the definition of cones that we introduce, effectively restricting to smaller cones of functions (see Definition 4.4), to estimate the diameters (see Proposition 4.6). This extra parameter takes care of the required global bounds on oscillations. Moreover, to ensure that the operator is a strict contraction on these new cones, we use a different operator and utilize additional techniques (see Proposition 4.5). Finally, we use the convergence rate of the new operator to obtain the desired bounds for the convergence rate of the normalized Ruelle–Perron–Frobenius operator.
We need this new finer estimate for the convergence rate (Theorem 4.1) to establish the computability of the eigenfunction $u_{\varphi}$ of the Ruelle–Perron–Frobenius operator. Furthermore, in Subsection 4.2, we deduce the computability of the Jacobian of the equilibrium state whose explicit expression is given by
$J_{\varphi}(x)\coloneqq\frac{u_{\varphi}(T(x))}{u_{\varphi}(x)}\exp(P(T,\varphi)-\varphi(x))\quad\text{for each }x\in X.$
@@
In Subsection 4.3, we establish the lower-computable openness of some subsets $U$ and $W$ of $\mathcal{P}(X)$ by expressing them as unions of uniformly computable sequences of lower-computable open sets. It is worth noting that the lower-computable openness of $U$ depends on the computability of $T$, while the lower-computable openness of $W$ depends on the computability of the Jacobian function $J_{\varphi}(x)$ in addition to the computability of $T$. In Subsection 4.4, we complete the proof of Theorem 1.1 and apply it to establish Theorem B. Here is a detailed version of Theorem B.
@@
There exists an algorithm with the following property:
For each septet $(X,\,D,\,\rho,\,\mathcal{S},\,\mathcal{S}^{\prime},\,\varphi,\,T)$ satisfying the Assumption B in Section 1, this algorithm outputs $\mu_{n}\in\mathcal{P}(X)$ supported on finitely many points of $\mathcal{S}$ such that the Wasserstein–Kantorovich distance
@@
where the measure $\mu$ is the unique equilibrium state for the restriction $T|_{D}$ and $\varphi$, after inputting the following data:
@@
    two rational constants $a_{0}>0$ and $v_{0}\in(0,1]$ with $\varphi\in C^{0,v_{0}}(X)$ and $a_{0}\geqslant\lvert\varphi\rvert_{v_{0},\rho}$,
@@
    three rational constants $\eta>0$, $\lambda>1,$ and $\xi>0$ satisfying that $T$ is distance-expanding with constants $\eta,\,\lambda$, and $\xi$ (in the sense of Definition 2.1),
@@
Next, in Section 5, we provide the proof of Theorem C. By Theorem 1.2, it suffices to check that the septet $\bigl{(}\widehat{\mathbb{C}},\,\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w),\,t\varphi_{f},\,f\bigr{)}$ satisfies the Assumption B in Section 1, where $\mathbb{Q}^{2}$ is defined as $\{a+bi:a,\,b\in\mathbb{Q}\}$, and $w$ is a repelling periodic point of the rational map $f$. The main difficulty here is to show that $\mathcal{J}_{f}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{f},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}f^{-n}(w)\bigr{)}$. We use the computability of the hyperbolic Julia sets proved by Braverman [Br04] to establish this statement.
All of our analyses for Theorems 1.1 and 1.2 should still work after necessary modifications if we replace the algorithms in the inputs by oracles. For the simplicity of the presentation, we leave the details to the reader.
Structure of the paper.
In Section 2, we introduce our notations and review some basic notions and results in ergodic theory and thermodynamic formalism in particular. Section 3 covers some basics of Computable Analysis. Section 4 is dedicated to the proofs of Theorems 1.1 and 1.2. In Subsection 4.1, we prove that the convergence of the iterations of a normalized Ruelle–Perron–Frobenius operator is at an exponential rate with explicit expressions for the related constants. Based on this, we demonstrate the computability of topological pressure $P(T,\varphi)$ and the function $u_{\varphi}$, which leads to the computability of the Jacobian of the equilibrium state in Subsection 4.2. Then, we prove Lemma 4.13 in Subsection 4.3 and establish Theorems 1.1 and 1.2 in Subsection 4.4. In Section 5, we introduce some results on the dynamical and algorithmic aspects of hyperbolic rational maps and prove Theorem C. In Section 6, we establish Theorem D, from which we derive Theorem E. In Section 7, we investigate and construct some computable systems whose equilibrium states are all non-computable, establishing Theorems F and G. In Subsection 7.1, we recall the computable map $T\colon\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ given in [GHR11, Subsection 4.1] and show that it satisfies the statements in Theorem F. Finally, in Subsection 7.2, we construct $\widehat{T}\coloneqq T\times T_{d}$ to establish Theorem G, where $T_{d}(x)\coloneqq dx\pmod{1}$ for $d\in\mathbb{N}$ and $x\in\mathbb{S}^{1}$.
@@ Acknowledgments
The authors would like to thank the anonymous referees for their valuable suggestions. I. Binder was partially supported by an NSERC Discovery grant. Z. Li and Q. He were partially supported by NSFC Nos. 12471083, 12101017, 12090010, 12090015, and BJNSF No. 1214021. Q. He was also partially supported by Peking University Funding Nos. 7101303303 and 6201001846. Y. Zhang was partially supported by NSFC Nos. 12161141002 and 12271432, and USTC–AUST Math Basic Discipline Research Center.
@@ 2. Preliminaries on thermodynamic formalism
In this section, we go over some notations, key concepts, and useful results in ergodic theory and thermodynamic formalism.
@@ 2.1. Notations
Denote $\mathbb{N}\coloneqq\{1,\,2,\,3,\,\ldots\}$ and $\mathbb{N}_{0}\coloneqq\{0\}\cup\mathbb{N}$. Set $\mathbb{N}^{*}\coloneqq\bigcup_{k\in\mathbb{N}}\mathbb{N}^{k}$.
For a subset $A\subseteq X$, the characteristic function of $A$ is denoted by $\mathbbm{1}_{A}$. Define the global characteristic function $\mathbbm{1}\coloneqq\mathbbm{1}_{X}$.
Let $(X,\rho)$ be a compact metric space. Let $C(X)$ and $C^{*}(X)$ denote the Banach space of continuous functions from $X$ to $\mathbb{R}$ equipped with the norm $\|u\|_{\infty}\coloneqq\sup\{\lvert u(x)\rvert:x\in X\}$ and the dual space of $C(X)$, respectively. The function $\varphi\colon X\rightarrow\mathbb{R}$ is said to be Hölder continuous (with respect to the metric $\rho$) with an exponent $\alpha\in(0,1]$ if there exists $C>0$ satisfying $\lvert\varphi(x)-\varphi(y)\rvert\leqslant C\rho(x,y)^{\alpha}$ for each pair of $x,\,y\in X$. Denote by $C^{0,\alpha}(X,\rho)$ the Banach space of real-valued Hölder continuous functions with an exponent $\alpha$. For each $\varphi\in C^{0,\alpha}(X,\rho)$, we define
$\lvert\varphi\rvert_{\alpha,\rho}\coloneqq\sup\bigl{\{}\lvert\varphi(x)-\varphi(y)\rvert/\rho(x,y)^{\alpha}:x,\,y\in X,\,x\neq y\bigr{\}}.$
More specifically, we say that a function $\phi\in C(X)$ is $c$-Lipschitz if $\phi\in C^{0,1}(X,\rho)$ satisfies that $\lvert\phi\rvert_{1,\rho}\leqslant c$. Denote by $c\operatorname{-Lip}(X)$ the set of real-valued $c$-Lipschitz continuous functions. We denote by $B_{\rho}(x,r)$ the open ball of radius $r>0$ centered at $x\in X$ and often omit the metric $\rho$ in the subscript when it is clear from the context.
Let $\mathcal{P}(X)$ and $\mathcal{M}(X,T)$ denote the set of Borel probability measures endowed with the weak topology and the set of $T$-invariant Borel probability measures on $X$, respectively. The Wasserstein–Kantorovich metric $W_{\rho}$ on $\mathcal{P}(X)$ metrizes the weak topology when $X$ is compact and is defined by
(2.1) $W_{\rho}(\mu,\nu)\coloneqq\sup\biggl{\{}\biggl{\lvert}\int\!f\,\mathrm{d}\mu-\int\!f\,\mathrm{d}\nu\biggr{\rvert}:f\in 1\operatorname{-Lip}(X)\biggr{\}}\quad\text{ for each pair of }\mu,\,\nu\in\mathcal{P}(X).$
Let $T\colon X\rightarrow X$ be a continuous map. We define the non-wandering set $\Omega(T)$ for $T$ by
@@
Moreover, for all $\varphi\in C(X)$, $n\in\mathbb{N}$, and $x\in X$, we define
$S_{n}\varphi(x)\coloneqq\sum\limits_{k=0}^{n-1}\varphi\circ T^{k}(x),$
where $T^{k}$ denotes the $k$-th iteration of the map $T$.
@@ 2.2. Basic concepts in ergodic theory
We begin with a brief introduction to distance-expanding maps, topologically transitive maps, and topologically exact maps. See e.g., [PU10, Chapters 3 and 4] for more details.
Given two constants $\eta>0$ and $\lambda>1$, assume that $(X,\rho)$ is a compact metric space and $T\colon X\rightarrow X$ is a continuous open map satisfying
@@
Then there exists $\xi>0$ such that
(2.4) $B(T(x),\xi)\subseteq T(B(x,\eta))\quad\text{ for each }x\in X.$
By (2.3), $T$ is injective on $B(x,\xi)$. Hence, for each $x\in X$, we can define the branch $T_{x}^{-1}\colon B(T(x),\xi)\rightarrow B(x,\eta)$ of the inverse map. Moreover, for all $x\in X$, $n\in\mathbb{N}$, the composition $T_{x_{0}}^{-1}\circ T_{x_{1}}^{-1}\circ\cdots\circ T_{x_{n-1}}^{-1}\colon B(T^{n}(x),\xi)\rightarrow X$ is well-defined, and is denoted by $T_{x}^{-n}$, where $x_{j}=T^{j}(x)$ for each integer $0\leqslant j\leqslant n-1$. Moreover, we have (see e.g. [PU10, Section 4.1])
$T^{-n}(A)=\bigcup_{x\in T^{-n}(y)}T^{-n}_{x}(A)\quad\text{ for all }y\in X\text{ and }A\subseteq B(y,\xi),$
and
(2.5) $\rho\bigl{(}T_{x}^{-n}(y),T_{x}^{-n}(z)\bigr{)}\leqslant\lambda^{-n}\rho(y,z)\quad\text{ for all }n\in\mathbb{N},\,x\in X,\text{ and }y,\,z\in B\bigl{(}T^{n}(x),\xi\bigr{)}.$
@@
Let $(X,\rho)$ be a compact metric space. A continuous map $T\colon X\rightarrow X$ is said to be distance-expanding with respect to the metric $\rho$ if there exist constants $\eta>0$ and $\lambda>1$ satisfying (2.3). Moreover, if $T$ is open, then we say that such a map $T$ is distance-expanding (with respect to $\rho$) with constants $\eta,\,\lambda$, and $\xi$ if the constants $\eta,\,\lambda$, and $\xi$ satisfy both (2.3) and (2.4).
We recall the definition of topologically transitive maps and topologically exact maps here (see e.g. [PU10, Definitions 4.3.1, 4.3.3]).
Definition 2.2.
Let $X$ be a topological space. A continuous map $T\colon X\rightarrow X$ is called topologically transitive if, for each pair of non-empty open sets $U,\,V\subseteq X$, there exists $n\in\mathbb{N}$ with $T^{n}(U)\cap V\neq\emptyset$. As a stronger property, we say that a continuous map $T\colon X\rightarrow X$ is topologically exact if, for each non-empty open set $U\subseteq X$, there exists $n\in\mathbb{N}$ with $T^{n}(U)=X$.
Under the above conventions, given constants $\eta,\,\xi>0$, and $\lambda>1$, a list of assumptions applied in the rest of this section is listed below:
@@
    The map $T\colon X\rightarrow X$ is open, continuous, topologically transitive, and distance-expanding with respect to $\rho$ with constants $\eta$, $\lambda$, and $\xi$.
Next, we recall some definitions and results in ergodic theory. For more details, we refer to [PU10, Chapter 2].
Let $X$ be a topological space. A measurable partition $\mathcal{A}$ of $X$ is a cover $\mathcal{A}=\{A_{j}:j\in J\}$ of $X$ consisting of finitely or countably many mutually disjoint Borel sets. Unless otherwise stated, we assume that our partitions are all measurable. For each partition $\mathcal{A}$ of $X$ and $x\in X$, the unique element of $\mathcal{A}$ containing $x$ is denoted by $\mathcal{A}(x)$.
Let $\mathcal{A}=\{A_{j}:j\in J\}$ and $\mathcal{B}=\{B_{k}:k\in K\}$ be two covers of $X$, where $J$ and $K$ are the corresponding index sets. We say that $\mathcal{A}$ is a refinement of $\mathcal{B}$ if for each $j\in J$, there exists $k=k(j)\in K$ such that $A_{j}\subseteq B_{k}$. The common refinement $\mathcal{A}\vee\mathcal{B}$ of $\mathcal{A}$ and $\mathcal{B}$ is a cover defined as $\mathcal{A}\vee\mathcal{B}\coloneqq\{A_{j}\cap B_{k}:j\in J,\,k\in K\}.$ For each continuous map $T\colon X\rightarrow X$, by $T^{-1}(\mathcal{A})$ we denote the partition $\bigl{\{}T^{-1}(A_{j}):j\in J\bigr{\}}$, and for each $n\in\mathbb{N}$, write
@@
Let $(X,\rho)$ be a compact metric space, $T\colon X\rightarrow X$ be a continuous map, and $\varphi\colon X\rightarrow\mathbb{R}$ be a continuous function (which is often called a potential). Denote by $\mathcal{Y}$ the set of sequences $\{\mathcal{V}_{n}\}_{n\in\mathbb{N}}$ of finite open covers of $X$ satisfying that $\operatorname{mesh}(\mathcal{V}_{n})$ converges to zero as $n$ tends to $\infty$, where the mesh of a cover $\mathcal{V}$ of $X$ is defined as $\operatorname{mesh}(\mathcal{V})\coloneqq\sup\{\operatorname{diam}(V):V\in\mathcal{V}\}.$ Define the topological pressure as the supremum of the following limits taken over all sequences $\{\mathcal{V}_{n}\}\in\mathcal{Y}$: